Six people are running a race. How many arrangements of winner, second place, and third place are possible?
A 18
B 36
C 108
D 120

Answer by stanbon(75887)

Six people are running a race. How many arrangements of winner, second place, and third place are possible?
----------------
6P3=6!/[6-3]! = 6*5*4=120 arrangements
Cheers,
Stan H.